# Simultaneous extension of the polynomial Central Sets Theorem and the polynomial Milliken-Taylor theorem

### v1

@@ 2   Our proofs
In the following theorem, we explicitly deduce some ultrafilters witnessing the polynomial extension of [13, Theorem 17.31]. However, we don’t know if the converse part of this theorem is true or not.
@@
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\in A.$
@@
Let $\Phi:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be the map defined by $\Phi(x_{1},x_{2},\dots,x_{m})=\sum^{m}_{i=1}a_{i}x_{i}$. Clearly $\Phi$ is a continuous map, and so we may apply [13, Theorem 3.49]. Let $\widetilde{\Phi}:\beta(\mathbb{N}^{m})\rightarrow\beta\mathbb{N}$ be the continuous extension of $\Phi$. Let $p\in E(K(\beta\mathbb{N},+))$, and $\bigotimes_{i=1}^{m}p\in\beta(\mathbb{N}^{m})$. Now,
$\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}\right]=\widetilde{\Phi}\left[p-\lim_{x_{1}}\cdots p-\lim_{x_{m}}(x_{1},\dots,x_{m})\right]$
@@
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $h\left(a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{1}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P\left(\sum_{i\in\beta_{m}}\sum_{t\in H(a_{i})}f_{i}(t)\right)\right)\right)\in A.$
@@
3   Separation of polynomial Milliken-Taylor systems
@@
Let $m\in\mathbb{N}$, and $f,g:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be two polynomials of finite degree defined by $f(x_{1},\dots,x_{m})=\sum a_{i_{1}\cdots i_{dn}}x^{i_{1}}_{1}\cdots x^{i_{m}}_{m}$, and $g(x_{1},\dots,x_{m})=\sum b_{j_{1}\cdots j_{m}}x^{j_{1}}_{1}\cdots x^{j_{m}}_{m}$. We say $f$ and $g$ are equivalent, i.e. $f\approx g$ if and only if, there exists a rational $\alpha\in\mathbb{Q}$ such that,
@@
For example the polynomials $f(x,y,z)=x^{3}+2xy^{2}z+xz+z^{2}$, and $g(x,y,z)=5x^{3}+10xy^{2}z+5xz+5z^{2}$ are equivalent.
@@
The following theorem shows that if two polynomials are equivalent then their Milliken-Taylor image for some sequences lies in the same color after partitioning $\mathbb{N}$ into finitely many pieces.
Theorem 3.5 (Separating Polynomial Milliken-Taylor).
@@ Acknowledgement
The first named author is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.

### v3

@@ 2  Our Proofs
The following theorem is a variation of Theorem 2.2. Proof is similar to the proof of above one and hence we omit it.
Theorem 2.3.
  1. 1.
    Let $r,m\in\mathbb{N}$, and let $\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$. Then there exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for each $i\in\{1,2,\ldots,m\}$, we have $\max H_{i}(F)<\min H_{i}(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have
      $\left\{\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right),\ldots,\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right):\right$.
      $\left.\hskip 216.81ptf_{i}\in G_{i}\text{ for every }i\in\mathbb{N}\right\}\subset C_{j}$
      for some $j\in\{1,2,\ldots,r\}$.
  2. 2.
    Let $r,m\in\mathbb{N}$, and let $a_{1},\dots,a_{m}\in\mathbb{N}$. Then for any $\mathbb{N}=\bigcup^{r}_{i=1}D_{i}$, there exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that such that for every $P\in T,$
    1. (a)
      if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for each $i\in\{1,2,\ldots,m\}$, we have $\max H_{i}(F)<\min H_{i}(G)$, and
    2. (b)
      whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$ we have for every $f_{i}\in G_{i},$ and $P_{1},\ldots,P_{m}\in T,$
      $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in D_{j}$
      for some $j\in\{1,2,\ldots,r\}.$
In the following theorem, we explicitly deduce some ultrafilters witnessing the polynomial extension of [14, Theorem 17.31]. However, we don’t know if the converse part of this theorem is true or not.
@@
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\in A.$
@@
In this proof we do not use [2, Theorem 3.2], but we use simple concept of $p-\lim.$ Let $\Phi:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be the map defined by $\Phi(x_{1},x_{2},\dots,x_{m})=\sum^{m}_{i=1}a_{i}x_{i}$. Clearly, $\Phi$ is a continuous map, and so we may apply [14, Theorem 3.49]. Let $\widetilde{\Phi}:\beta(\mathbb{N}^{m})\rightarrow\beta\mathbb{N}$ be the continuous extension of $\Phi$. Let $p\in E(K(\beta\mathbb{N},+))$, and $\bigotimes_{i=1}^{m}p\in\beta(\mathbb{N}^{m})$. Now,
$\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}p\right]=\widetilde{\Phi}\left[p-\lim_{x_{1}}\cdots p-\lim_{x_{m}}(x_{1},\dots,x_{m})\right]$
@@
The following theorem is a variation of Theorem 2.4.
Theorem 2.5.
Let $m\in\mathbb{N},$ $a_{1},\dots,a_{m}\in\mathbb{N}$, $p_{1},\ldots,p_{m}\in E\left(K(\beta\mathbb{N},+)\right)$, and $A\in a_{1}p_{1}+\dots+a_{m}p_{m}$. Then for every $T\in\mathcal{P}_{f}(\mathbb{P})$, there exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for each $i\in\{1,2,\ldots,m\}$, we have $\max H_{i}(F)<\min H_{i}(G)$, and
  2. 2.
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\in A.$
In fact we may consider $p_{1},\ldots,p_{m}\in E\left(\mathcal{J}_{p},+\right).$
Proof.
The proof is similar to the proof of Theorem 2.4. Let $\Phi:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be the map defined by $\Phi(x_{1},x_{2},\dots,x_{m})=\sum^{m}_{i=1}a_{i}x_{i}$. Clearly, $\Phi$ is a continuous map, and so we may apply [14, Theorem 3.49]. Let $\widetilde{\Phi}:\beta(\mathbb{N}^{m})\rightarrow\beta\mathbb{N}$ be the continuous extension of $\Phi$. Let $p_{1},\ldots,p_{m}\in E(K(\beta\mathbb{N},+))$, and $\bigotimes_{i=1}^{m}p_{i}\in\beta(\mathbb{N}^{m})$. Now,
$\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}p_{i}\right]=\widetilde{\Phi}\left[p_{1}-\lim_{x_{1}}\cdots p_{m}-\lim_{x_{m}}(x_{1},\dots,x_{m})\right]$
$=p_{1}-\lim_{x_{1}}\cdots p_{m}-\lim_{x_{m}}\Phi(x_{1},\dots,x_{m})$
$=p_{1}-\lim_{x_{1}}\dots p_{m}-\lim_{x_{m}}\sum^{m}_{i=1}a_{i}x_{i}$
$=a_{1}\cdot p_{1}+\dots+a_{m}\cdot p_{m}.$
Let $A\in a_{1}p_{1}+\dots+a_{m}p_{m}=\widetilde{\Phi}\left[\bigotimes_{i=1}^{m}p_{i}\right]$. Hence there exists $B\in\bigotimes_{i=1}^{m}p_{i}$ such that $\Phi[B]\subseteq A$. Now from Corollary 1.30, our result follows.
@@
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $h\left(a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H(G_{i})}f_{i}(t)\right)\right)\right)\in A.$
@@
The following corollary is an immediate variation of Corollary 2.6.
Corollary 2.7.
Let $m\in\mathbb{N},$ $h:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be any polynomial with no constant term. Let $a_{1},\dots,a_{m}\in\mathbb{N}$, $p_{1},\ldots,p_{m}\in E\left(K(\beta\mathbb{N},+)\right)$, and $A\in\tilde{h}(a_{1}p_{1}+\dots+a_{m}p_{m})$. Then for every $T\in\mathcal{P}_{f}(\mathbb{P})$, there exists functions $\alpha_{1},\ldots,\alpha_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H_{1},\ldots,H_{m}:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
  1. 1.
    if $F,G\in\mathcal{P}_{f}(\tau)$ and $F\subsetneq G$ then for every $i\in\{1,\ldots,m\}$, we have $\max H_{i}(F)<\min H_{i}(G)$, and
  2. 2.
    whenever $(G_{i})_{i}\in\mathcal{P}_{f}(\tau)$, is a sequence of functions, i.e. $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}\subsetneq\cdots$, then for every $P_{1},\ldots,P_{m}\in T,$ $\beta_{1}<\cdots<\beta_{m}$ in $\mathcal{P}_{f}(\mathbb{N})$, we have
    $h\left(a_{1}\cdot\left(\sum_{i\in\beta_{1}}\alpha_{1}(G_{i})+P_{1}\left(\sum_{i\in\beta_{1}}\sum_{t\in H_{1}(G_{i})}f_{i}(t)\right)\right)+\cdots+a_{m}\cdot\left(\sum_{i\in\beta_{m}}\alpha_{m}(G_{i})+P_{m}\left(\sum_{i\in\beta_{m}}\sum_{t\in H_{m}(G_{i})}f_{i}(t)\right)\right)\right)\in A.$
In fact we may consider $p_{1},\ldots,p_{m}\in E\left(\mathcal{J}_{p},+\right).$
Proof.
Let $A\in\tilde{h}(a_{1}p_{1}+\dots+a_{m}p_{m})$, and let $B\in a_{1}p_{1}+\dots+a_{m}p_{m}$ be such that $h[B]\subseteq A.$ Now Theorem 2.5 directly implies our result. ∎
Though we don’t know the ultrafilter characterization of the Theorems 2.4, 2.5, but we can completely characterize them for the set $T=\{i\}$, where $i(x)=x$ for all $x\in\mathbb{N}.$ The following theorem is a variation of [14, Theorem 17.31] for idempotents of $(\mathcal{J},+)$.
Theorem 2.8.
Let $T=\{i\}\in\mathcal{P}_{f}(\mathbb{P})$, and $A\subseteq\mathbb{N}$.
  1. 1.
    if AA satisfies the conclusion of Theorem 2.4 if and only if there exists idempotent ultrafilter $p\in(\mathcal{J},+)$ such that $A\in a_{1}p+\cdot+a_{m}p.$
  2. 2.
    if AA satisfies the conclusion of Theorem 2.5 if and only if there exists idempotent ultrafilters $p_{1},\ldots,p_{m}\in(\mathcal{J},+)$ such that $A\in a_{1}p_{1}+\cdot+a_{m}p_{m}.$
Proof.
The proof of the sufficient condition of both $(1)$ and $(2)$ is similar to the proof of Theorem 2.4, 2.5. For the necessary part, both proofs are similar. So we prove the necessary part of $(2).$
Let $B\subseteq\mathbb{N}^{m}$ be such that $(x_{1},\ldots,x_{m})\in B\text{ if }\sum_{i=1}^{m}a_{i}x_{i}\in A.$ Then the set $B$ satisfies the combinatorial conclusion of the Corollary 1.26. Hence there exist idempotents $p_{1},\ldots,p_{m}$ in $(\mathcal{J},+)$, such that $B\in\bigotimes_{i=1}^{m}p_{i}.$ Letting $\Phi:\mathbb{N}^{m}\rightarrow\mathbb{N}$ be the map defined by $\Phi(x_{1},x_{2},\dots,x_{m})=\sum^{m}_{i=1}a_{i}x_{i}$, from Theorem 2.5, we have
$\Phi[B]\in\tilde{\Phi}\left(\bigotimes_{i=1}^{m}p_{i}\right)=a_{1}p_{1}+\cdots+a_{m}p_{m}.$
But from the construction $\Phi[B]\subseteq A.$ Hence $A\in a_{1}p_{1}+\cdots+a_{m}p_{m}.$ This completes our proof. ∎
3  Separation of the Polynomial Milliken-Taylor Systems
@@
Let $m,n\in\mathbb{N}$, and $f:\mathbb{N}^{m}\rightarrow\mathbb{N}$, $g:\mathbb{N}^{n}\rightarrow\mathbb{N}$ be two polynomials of finite degree defined by $f(x_{1},\dots,x_{m})=\sum a_{i_{1}\cdots i_{m}}x^{i_{1}}_{1}\cdots x^{i_{m}}_{m}$, and $g(x_{1},\dots,x_{n})=\sum b_{j_{1}\cdots j_{n}}x^{j_{1}}_{1}\cdots x^{j_{n}}_{n}$. We say $f$ and $g$ are equivalent, i.e. $f\approx g$ if and only if,
  1. 1.
    both $f,g$ have same degree, and
  2. 2.
    there exists a rational $\alpha\in\mathbb{Q}$ such that,
@@
For example the polynomials $f(x,y,z)=x^{3}+2xy^{2}z+xz+z^{2}$, and $g(x,y,z)=5^{3}x^{3}+12\cdot 5^{4}xy^{2}z+5^{2}xz+5^{2}z^{2}$ are equivalent.
@@
The following theorem shows that if two polynomials are equivalent then their Milliken-Taylor image for some sequences lie in the same color after partitioning $\mathbb{N}$ into finitely many pieces.
Theorem 3.5 (Separating Polynomial Milliken-Taylor theorem).
@@ Acknowledgement
The first named author is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927. We are also thankful to Amit Roy for his technical assistantship.